7. The ABC company estimates that its employees are currently averaging $100 or less per month in tax-deferred savings. A sample of 50 employees will be used to test ABC's hypothesis abo Assume that the employee tax-deferred savings amounts have a standard deviation of $80 and that a 5% level of significance will be used for the test. What is the probability of making a Type II error if the "True" monthly savings average $125? What is the probability of making a Type I error? ut the current level of savings activity among the population of its employees. 
Answer #1

Ans:

z=1.645

sample mean cut off value=100+1.645*(80/sqrt(50))=118.61

when true mean is 125:

z=(118.61-125)/(80/sqrt(50))

z=-0.565

P(Type II error)=P(z<-0.565)=0.2860